# A Comprehensive Survey on Vector Database: Storage and Retrieval Technique, Challenge

### v2

@@ Abstract
Vector databases (VDBs) have emerged to manage high-dimensional data that exceed the capabilities of traditional database management systems, and are now tightly integrated with large language models as well as widely applied in modern artificial intelligence systems. Although relatively few studies describe existing or introduce new vector database architectures, the core technologies underlying VDBs, such as approximate nearest neighbor search, have been extensively studied and are well documented in the literature. In this work, we present a comprehensive review of the relevant algorithms to provide a general understanding of this booming research area. Specifically, we first provide a review of storage and retrieval techniques in VDBs, with detailed design principles and technological evolution. Then, we conduct an in-depth comparison of several advanced VDB solutions with their strengths, limitations, and typical application scenarios. Finally, we also outline emerging opportunities for coupling VDBs with large language models, including open research problems and trends, such as novel indexing strategies. This survey aims to serve as a practical resource, enabling readers to quickly gain an overall understanding of the current knowledge landscape in this rapidly developing area.
Index Terms:
Vector Database, Retrieval, Storage, Large Language Models.
I Introduction
@@
Vectors, particularly those in high-dimensional spaces, are mathematical representations of data, encoding the semantic and contextual information of entities such as text, images, audio, and video[1, 2]. These vectors are generally generated through some related machine learning models, and the generated vectors are usually high-dimensional and can be used for similarity comparison. The step of converting original unstructured data into vectors is the foundation of many artificial intelligence (AI) applications (including large language models (LLMs)[3], question-answering systems[4, 5], image recognition[6], recommendation systems[7, 8], etc.). However, in terms of managing and retrieving high-dimensional vector data, traditional databases designed for handling structured data are often inadequate. Vector databases, on the other hand, provide a specialized solution to these challenges.
Vector Databases (VDBs) are tools specifically designed to efficiently store and manage high-dimensional vectors. Specifically, VDBs store information as high-dimensional vectors, which are mathematical representations of data features or attributes[9]. Depending on the complexity and granularity of the underlying data, the dimensions of these high-dimensional vectors usually range from dozens to thousands. Unlike traditional relational databases, VDBs provide efficient mechanisms for large-scale storage, management, and search of high-dimensional vectors[10, 11, 12]. These mechanisms bring various efficient functions to VDBs, such as supporting semantic similarity search, efficiently managing large-scale data, and providing low-latency responses. These functions make VDBs increasingly integrated into AI-based applications.
VDBs have two core functions: vector storage and vector retrieval. The vector storage function relies on techniques such as quantization, compression, and distributed storage mechanisms to improve efficiency and scalability. The retrieval function of VDBs relies on specialized indexing techniques, including tree-based methods, hashing methods [13], graph-based models, and quantization-based techniques[14]. These indexing techniques optimize high-dimensional similarity search by reducing computational cost and improving search performance. In addition, hardware acceleration and cloud-based technologies have further enhanced the capabilities of VDBs, making them suitable for large-scale and real-time applications[15, 16, 17].
Therefore, compared with traditional databases, vector databases (VDBs) have three significant advantages: (1) Vector databases possess efficient and accurate vector retrieval capabilities. The core function of VDBs is to retrieve relevant vectors through vector similarity (distance), and this function is also the core of applications such as natural language processing (NLP), computer vision, and recommendation systems[18, 19]. In contrast, traditional databases can only query data based on exact matching or predefined conditions, and this kind of query method is relatively slow and often does not consider the semantic information of the data itself. (2) Vector databases support the storage and query of complex and unstructured data. VDBs can store and search data with high complexity and granularity, such as text, images, audio, video, etc. However, traditional databases, such as relational databases, are difficult to store this kind of data information well[20]. (3) Vector databases have high scalability and real-time processing capabilities. Traditional databases face challenges in effectively handling unstructured data and large volumes of real-time datasets[21]. However, VDBs can process vector data at large scale and in real time, which is crucial for modern data science and artificial intelligence applications[22]. By using technologies such as sharding[23], partitioning, caching, and replication, VDBs can distribute workloads and optimize resource utilization across multiple machines or clusters. Traditional databases, on the other hand, may face scalability bottlenecks, latency issues, or concurrency conflicts when handling big data[19].
@@
Recent surveys on VDBs primarily cover fundamental concepts and practical applications of VDBs and vector database management systems. Some studies [12, 24, 19] focus on the workflow and technical challenges of VDBs, including key aspects such as query processing, optimization, and execution techniques. And some works [25] explore the critical role of VDBs in modern generative AI applications and provide an outlook on the future of VDBs. While these studies have their respective focuses, they do not provide a comprehensive survey of the overall storage and search technologies in VDBs, nor deliver a thorough analysis comparing the capabilities of existing VDBs. Furthermore, there is limited exploration of how these systems can integrate with rapidly advancing AI technologies, such as large language models (LLMs), to support modern data-intensive applications.
This gap shows the need for a comprehensive survey to consolidate current knowledge and uncover key research challenges related to VDB. To address this, our survey makes the following core contributions:
  • We systematically review storage and retrieval techniques in VDBs, outlining their design principles and technological evolution.
  • We provide a detailed comparative analysis of existing VDB solutions, highlighting their strengths, limitations, and typical application scenarios.
  • We synthesize the main challenges, recent advancements, and future directions for VDBs, including open research problems and trends such as novel indexing strategies, adaptive query optimization, and integration with advanced AI capabilities.
@@
This paper comprehensively summarizes the technologies related to vector databases, and systematically tests the performance of existing open-source vector databases. It also provides an outlook on the challenges that vector databases will face in the future. Through the summary of this paper, researchers can deepen their understanding of the field of vector databases. Figure  1 shows the overall framework of the paper, and we also construct a classification system of storage and search technologies for VDBs, as shown in Figure  2.
II Storage
@@
Efficient data management strategies are essential for the performance and scalability of VDBs. This section explores four key techniques: sharding and partitioning for data distribution, caching for reducing query latency, and replication for ensuring availability and fault-tolerance.
II-A Sharding
Sharding is a technique that distributes a VDB across multiple machines or clusters, called shards, based on specific criteria, such as a hash function or a key range. It is a foundational strategy for achieving scalability, load balancing, and fault tolerance in distributed systems by dividing the dataset into smaller, manageable units. Different sharding methods are designed to address specific challenges, such as balancing data distribution, minimizing hotspots, reducing query latency, and supporting dynamic scaling. This section covers three widely adopted strategies: range-based sharding, hash-based sharding, and geographic sharding.
@@ II-A Sharding
Range-Based Sharding. The most straightforward sharding method is range-based sharding, which partitions vector data across shards by dividing it into non-overlapping key intervals, or ranges, based on sorted keys[26]. For example, users can shard a VDB by dividing the ID column of the vector data into different ranges, such as 0-999, 1000-1999, 2000-2999, and so on, with each range assigned to a specific shard. Each range corresponds to a shard. This way, users can query the vector data more efficiently by specifying the shard name or range. Range-based sharding is simple to implement and provides efficient query performance for range-based queries, such as retrieving all vectors within a specific time period or ID range. However, this method often leads to data skew and uneven load distribution if key values are not uniformly distributed. Range-based sharding is particularly suitable for use cases where data access patterns are predictable, such as time-series analytics or sequential ID-based queries.
Figure 2: Taxonomy of Vector Database (VDB) Storage and Search Technologies.
@@
Hash-Based Sharding. Another common sharding method in VDBs is hash-based sharding[27, 28, 26], which assigns vector data into different shards based on the hash value of a key column or a set of columns. For example, users can shard a VDB by applying a hash function to the ID column of the vector data. For example, users can distribute the vector data evenly across the shards and avoid hotspots. However, traditional hash-based methods, such as modulo hashing, introduce challenges during cluster scaling, suffering from significant data redistribution when adding or removing shards. To address this, consistent hashing is often employed in distributed VDBs. Unlike traditional hashing, consistent hashing maps both data and nodes to a ring-like hash space. This ensures that only a small fraction of data (O(1/N)) is remapped when nodes are added or removed. Consistent hashing minimizes reorganization overhead during cluster scaling while maintaining balanced data distribution. Additionally, virtual nodes (replicas of physical nodes on the hash ring) further enhance load balancing, mitigating hotspots. Although consistent hashing introduces slight computational complexity, its advantages in dynamic environments make it a preferred choice for modern distributed systems[29, 30, 31]
Geographic Sharding. Another approach is geographic sharding (also known as geo-sharding), which distributes vector data across shards based on geographic attributes [32, 33], such as user region or location. For instance, a VDB can be partitioned by a “region” field, assigning each shard to a specific geographic zone (e.g., “North America,” “Europe,” or “Asia”). This method is particularly useful for latency-sensitive localized queries, such as recommendation systems or geospatial analytics. By storing data in shards physically closer to users, geo-sharding reduces cross-region network latency and helps comply with data sovereignty regulations. Additionally, combined with dynamic load balancing, it can adapt to shifting geographic access patterns for optimal performance.
II-B Partitioning
Partitioning typically refers to dividing data within a single database instance into multiple logical subsets based on rules such as range, list, and K-Means, where all partitioned data remains within the same physical system. This approach improves query efficiency, facilitates parallel processing, and optimizes resource usage by organizing data into logical groups. For instance, users can partition a VDB by color values such as red, yellow, and blue, where each partition contains vector data associated with a specific “color” column value. Queries can then target specific partitions, avoiding unnecessary scans of unrelated data and improving performance. While sharding focuses on distributing data across multiple machines in a distributed environment, partitioning primarily organizes data within a single machine or node. These two techniques serve different purposes but can be used together synergistically. In practical applications, sharding achieves horizontal scalability by distributing data across nodes, while partitioning optimizes local data organization within each node. By combining the two, systems can balance global scalability and localized query performance—sharding handles load distribution for large datasets, and partitioning ensures efficient data access within individual shards. Partitioning strategies vary depending on the characteristics of the data and application requirements. This section discusses four common partitioning methods: range partitioning, list partitioning, k-means partitioning, and hash-based Partitioning.
@@ II-B Partitioning
Range-based Partitioning. Range-based partitioning is a method widely used in VDBs, where data is divided into non-overlapping key ranges to form partitions[34, 35]. Each range corresponds to a specific subset of data based on a sorted key (e.g., timestamps, numeric IDs). Similar to the strength of range-based sharding, range-based partitioning is particularly efficient for range-based queries, as it allows the system to target specific partitions directly. For example, users can partition a VDB by date ranges, such as monthly or quarterly. This way, users can query the vector data more efficiently by specifying the partition name or range.
@@
List-based Partitioning. Another way that partitioning works in VDB is by using a list partitioning method, which assigns vector data to different partitions based on their value lists of a key column or a set of columns[36, 37]. For example, users can partition a vector database by color values, such as red, yellow, and blue. Each partition contains vector data that have a given value of the color column. This way, users can query the vector data more easily by specifying the partition name or list.
@@
K-Means Partitioning. A third partitioning method for VDBs is k-means partitioning[20], which divides vector data into a predetermined number (k) of clusters. Each cluster represents a partition, with vectors within the same cluster being similar to each other and vectors between clusters differing significantly. Similar vectors are placed in the same partition, which improves query efficiency. However, for large-scale datasets, the computational cost of k-means clustering can be high, especially when frequent updates necessitate re-clustering.
@@
Hash-based Partitioning. Alternatively, some VDBs adopt hash-based partitioning, such as consistent or uniform hashing [38, 39, 40]. Specifically, the hashing partitioning strategy uses a hash function to map data to different partitions. The hash value of each data point determines which partition it belongs to. In VDBs, the hash value is typically calculated based on certain features of the vector (for example, values of specific dimensions or the entire vector). The hash function can evenly distribute data across partitions, preventing any single partition from storing too much data. However, when the data distribution is uneven, it may lead to some partitions becoming overloaded. Additionally, when the node number changes, hash-based partitioning may require the redistribution of the large dataset, which can incur significant overhead.
II-C Caching
Caching is a technique that stores frequently accessed or recently used data in a fast and accessible memory, such as RAM, to reduce latency and improve data retrieval performance. Caching can be used in VDBs to speed up similarity search and vector retrieval. In traditional database systems, utilizing in-memory key-value stores (e.g., Redis) as a caching layer represents a widely adopted approach. This methodology generates cache keys by combining query parameters (such as specific column values for matching). When a target key is found in the cache, Redis serves the corresponding data; otherwise, a cache miss occurs, necessitating database queries to fetch the required information. However, this key-value paradigm proves unsuitable for VDBs, as they store high-dimensional vector data rather than structured data, making it virtually impossible to reuse identical vectors across queries. To address these challenges, it is important to explore general caching algorithms and their applicability to VDBs. This section discusses four common caching methods: first-in first-out (FIFO), least recently used (LRU), most recently used (MRU), and least frequently used (LFU).
First-In First-Out (FIFO). FIFO algorithm is a fundamental cache eviction strategy that operates on the principle of “first in, first out”: when the cache space is exhausted, it prioritizes removing the earliest stored vector data [41, 42]. This algorithm maintains a simple queue structure: new data is always appended to the tail, while eviction removes the oldest data from the head[43]. FIFO is particularly suitable for vector data scenarios with stable access patterns and no distinct hotspots, such as time-series vector data collected at fixed intervals. While its implementation is simple and efficient (with O(1) time complexity), it may inadvertently evict frequently used data due to disregarding access frequency. A typical application includes real-time processing systems for industrial sensor data, where the timeliness of historical vector data often outweighs its reuse value.
Least Recently Used (LRU). One commonly employed caching strategy in VDBs is the least recently used (LRU) policy, which evicts the least recently accessed vector data when the cache reaches its capacity. This approach ensures that the cache retains the most relevant or frequently queried vectors, thereby improving the likelihood of cache hits [43]. For example, Redis, a widely used in-memory database, implements LRU caching to manage vector data and facilitate efficient vector similarity search. However, LRU may not effectively retain data with long-term or periodic popularity, as it only considers recent access history. LRU is particularly suitable for scenarios with strong temporal locality, such as recommendation systems and real-time search applications, where recently accessed vectors are highly likely to be queried again.
Most Recently Used (MRU). The Most Recently Used (MRU)[44] algorithm is another caching strategy employed in VDBs that prioritizes evicting the most recently accessed vector data when cache capacity is reached. This approach operates under the assumption that recently queried vectors are less likely to be accessed again in the short term, thereby retaining less recently used data that may be needed in the future. MRU has been adopted in certain storage systems and applications with transient or one-time access patterns, such as data streaming or processing workloads, where immediate reuse of recently accessed vectors is unlikely. However, MRU is suboptimal for workloads with strong temporal locality, as it may prematurely evict data that will soon be accessed again.

### v3

@@ Abstract
[Objective] As high‑dimensional vector data increasingly surpasses the processing capabilities of traditional database management systems, Vector Databases (VDBs) have emerged and become tightly integrated with large language models, being widely applied in modern artificial intelligence systems. However, existing research has primarily focused on underlying technologies such as approximate nearest neighbor search, with relatively few studies providing a systematic architectural‑level review of VDBs or analyzing how these core technologies collectively support the overall capacity of VDBs. This survey aims to offer a comprehensive overview of the core designs and algorithms of VDBs, establishing a holistic understanding of this rapidly evolving field. [Methods] First, we systematically review the key technologies and design principles of VDBs from the two core dimensions of storage and retrieval, tracing their technological evolution. Next, we conduct an in‑depth comparison of several mainstream VDB architectures, summarizing their strengths, limitations, and typical application scenarios. Finally, we explore emerging directions for integrating VDBs with large language models, including open research challenges and trends such as novel indexing strategies. [Conclusions]This survey serves as a systematic reference guide for researchers and practitioners, helping readers quickly grasp the technological landscape and development trends in the field of vector databases, and promoting further innovation in both theoretical and applied aspects.
keywords:
[…]
1 Introduction
@@
Vectors, particularly those in high-dimensional spaces, are mathematical representations of data, encoding the semantic and contextual information of entities such as text, images, audio, and video [cao2024knowledge, 10.1145/3150226]. These vectors are generally generated through some related machine learning models, and the generated vectors are usually high-dimensional and can be used for similarity comparison. The step of converting original unstructured data into vectors is the foundation of many artificial intelligence (AI) applications (including large language models (LLMs) [zhao2023survey], question-answering systems [allam2012question, biancofiore2024interactive], image recognition [7445232], recommendation systems [zhao2024recommender, liu2024multimodal], etc.). However, in terms of managing and retrieving high-dimensional vector data, traditional databases designed for handling structured data are often inadequate. Vector databases (VDBs), on the other hand, provide a specialized solution to these challenges.
@@
VDBs are tools specifically designed to efficiently store and manage high-dimensional vectors. Specifically, VDBs store information as high-dimensional vectors, which are mathematical representations of data features or attributes [10.1145/3382080]. Depending on the complexity and granularity of the underlying data, the dimensions of these high-dimensional vectors usually range from dozens to thousands. Unlike traditional relational databases, VDBs provide efficient mechanisms for large-scale storage, management, and search of high-dimensional vectors [wang2024efficient, kraska2018case, xie2023bri]. These mechanisms bring various efficient functions to VDBs, such as supporting semantic similarity search, efficiently managing large-scale data, and providing low-latency responses. These functions make VDBs increasingly integrated into AI-based applications. VDBs have two core functions: vector storage and vector retrieval. The vector storage function relies on techniques such as quantization, compression, and distributed storage mechanisms to improve efficiency and scalability. The retrieval function of VDBs relies on specialized indexing techniques, including tree-based methods, hashing methods [zhao2023towards], graph-based models, and quantization-based techniques [wang2023graph]. These indexing techniques optimize high-dimensional similarity search by reducing computational cost and improving search performance. In addition, hardware acceleration and cloud-based technologies have further enhanced the capabilities of VDBs, making them suitable for large-scale and real-time applications [li2019approximate, karthik2024bang, jegou2022faiss].
Consequently, VDBs offer three key benefits over traditional databases: (1) VDBs can retrieve vectors accurately and efficiently.VDBs’ primary purpose is to retrieve pertinent vectors based on vector similarity (distance); this function is also at the heart of applications like recommendation systems, computer vision, and natural language processing (NLP) [mohoney2023high, pan2024vector]. In contrast, traditional databases can only query data based on exact matching or predefined conditions, and this kind of query method is relatively slow and often does not consider the semantic information of the data itself. (2) VDBs can be used to store and query complex and unstructured data. Text, images, audio, video, and other types of data can all be stored and searched with great granularity and complexity using VDBs. However, traditional databases, such as relational databases, are difficult to store this kind of data information well [pan2024survey]. (3) VDBs are very scalable and have real-time data processing capabilities. Large volumes of real-time datasets and unstructured data are difficult for traditional databases to handle efficiently [rao2019big].
However, due to their ability to process vector data on a large scale and in real time, VDBs are crucial for modern data science and artificial intelligence [wang2024starling]. Through the use of technologies like caching, replication, partitioning, and sharding [su2024vexless], VDBs can optimize resource utilization and divide workloads among multiple machines or clusters. However, when dealing with big data, traditional databases might experience concurrency conflicts, latency problems, or scalability bottlenecks [pan2024vector].
@@
Recent surveys on VDBs primarily cover fundamental concepts and practical applications of VDBs and vector database management systems. Some studies [xie2023bri, taipalus2024vector, pan2024vector] focus on the workflow and technical challenges of VDBs, including key aspects such as query processing, optimization, and execution techniques. And some works [joshiintroduction] explore the critical role of VDBs in modern generative AI applications and provide an outlook on the future of VDBs. While these studies have their respective focuses, they do not provide a comprehensive survey of the overall storage and search technologies in VDBs, nor deliver a thorough analysis comparing the capabilities of existing VDBs. Furthermore, there is limited exploration of how these systems can integrate with rapidly advancing AI technologies, such as large language models (LLMs), to support modern data-intensive applications.
This research gap emphasizes the need for a thorough survey that aims to integrate the existing body of knowledge regarding VDBs and identify the main research issues that require immediate attention. The following are the main contributions our survey makes to address this issue:
  • We organized and reviewed the storage and retrieval technologies in VDBs in a methodical manner.
  • To highlight the advantages and disadvantages of each open-source VDB, we conducted thorough performance testing and comparison.
  • We went into great detail about the primary issues that VDBs are currently facing and how they integrate with Large Language Models (LLMs).
@@
This paper comprehensively summarizes the technologies related to VDBs and systematically tests the performance of existing open-source VDBs. It also provides an outlook on the challenges that VDBs will face in the future. Through the summary of this paper, researchers can deepen their understanding of the field of VDBs. Figure 1 shows the overall framework of the paper, and we also construct a classification system of storage and search technologies for VDBs, as shown in Figure 2.
Figure 2: Taxonomy of VDB Storage and Search Technologies.
2 Storage
@@
Efficient data management strategies are essential for the performance and scalability of VDBs. This section explores four key techniques: sharding and partitioning for data distribution, caching for reducing query latency, and replication for ensuring availability and fault tolerance.
2.1 Sharding
Sharding spreads a VDB across several computers or clusters, called shards, by routing records using a rule such as a hash or a key range. By cutting the dataset into smaller pieces, sharding helps the system scale, shares work evenly, and adds resilience when a single node fails. Different sharding methods target specific goals—such as easy expansion, hotspot reduction, fair size balancing, and quicker queries—so designers pick the one that fits their load. Here we review three popular schemes: range-based, hash-based, and geographic sharding.
@@ 2.1 Sharding
Range-Based Sharding. Range-based sharding is the simplest approach; it splits sorted keys into clean, non-overlapping intervals and sends each interval to its own shard  [oracle2021shard]. For example, users can shard a VDB by dividing the ID column of the vector data into different ranges, such as 0-999, 1000-1999, 2000-2999, and so on, with each range assigned to a specific shard. Each range corresponds to a shard. This way, users can query the vector data more efficiently by specifying the shard name or range. Range-based sharding provides efficient query performance and is simple to use for range-based queries, such as retrieving all vectors within a specified time frame or ID range. However, if key values are not distributed uniformly, this approach frequently results in data skew and uneven load distribution. Use cases like time-series analytics or sequential ID-based queries, where data access patterns are predictable, are best suited for range-based sharding.
@@
Hash-Based Sharding. Another common sharding method in VDBs is hash-based sharding [costa2015sharding, author2024ebook, oracle2021shard], which assigns vector data into different shards based on the hash value of a key column or a set of columns. For example, users can shard a VDB by applying a hash function to the ID column of the vector data. This lets users spread vector data evenly across the shards and avoid hotspots. But traditional hash-based methods, like modulo hashing, cause problems when scaling a cluster. When users add or remove shards, a large amount of data must be moved. To fix this, many distributed VDBs use consistent hashing. Unlike regular hashing, consistent hashing puts both data and nodes on a ring-shaped hash space. When a node is added or removed, only a small part of the data (O(1/N)) needs to move. This helps reduce the amount of data that must be changed during scaling. It also keeps the data balanced across the system. Also, virtual nodes (copies of physical nodes on the hash ring) help with load balancing. They spread the traffic more evenly and reduce hotspots. Although consistent hashing introduces slight computational complexity, its advantages in dynamic environments make it a preferred choice for modern distributed systems [mirrokni2018consistent, karger1999web, karger1997consistent].
Geographic Sharding. This method puts data into shards based on location, such as user region  [taft2020cockroachdb, zhang2013transaction]. A VDB can split data using a "region" field and send it to different zones like "North America" or "Europe." Geo-sharding helps with apps that need low delay, like recommendation systems. Storing data close to users lowers network delay. It also helps follow local data regulations.
2.2 Partitioning
Partitioning means splitting data in one database into smaller parts based on simple rules like range, list, or K-means. All the data stays in the same physical system. This method helps make queries faster. For example, users can split a VDB by color values like red, yellow, and blue. Each part holds vector data with a certain value in the “color” column. Queries can then target specific partitions, avoiding unnecessary scans of unrelated data. While sharding is used to distribute data across multiple machines, it helps systems scale horizontally. Partitioning, on the other hand, is used to organize data within a single machine to improve local access and performance. Using both together helps the VDBs handle large data and fast queries—sharding spreads the data across many machines, and partitioning makes data access faster inside each shard. Partitioning strategies change based on the type of data and what the application needs. This section looks at four common methods. These are range partitioning, list partitioning, k-means partitioning, and hash-based partitioning. Each method has its own use case and works best under different conditions.
@@ 2.2 Partitioning
Range-based Partitioning. Range-based partitioning is a method widely used in VDBs, where data is divided into non-overlapping key ranges to form partitions [dewitt1992parallel, dewitt1990hybrid]. Each range corresponds to a specific subset of data based on a sorted key (e.g., timestamps, numeric IDs). Similar to the strength of range-based sharding, range-based partitioning is particularly efficient for range-based queries, as it allows the system to target specific partitions directly. For example, users can partition a VDB by date ranges, such as monthly or quarterly. This way, users can query the vector data more efficiently by specifying the partition name or range.
@@
List-based Partitioning. Another way that partitioning works in VDB is by using a list partitioning method, which assigns vector data to different partitions based on their value lists of a key column or a set of columns [hobbs2011oracle, hotka2002oracle9i]. For example, users can partition a vector database by color values, such as red, yellow, and blue. Each partition contains vector data that has a given value of the color column. This way, users can query the vector data more easily by specifying the partition name or lists.
@@
K-Means Partitioning. A third partitioning method for VDBs is k-means partitioning [pan2024survey], which divides vector data into a predetermined number (k) of clusters. Each cluster represents a partition, with vectors within the same cluster being similar to each other and vectors between clusters differing significantly. Similar vectors are placed in the same partition, which improves query efficiency. However, for large-scale datasets, the computational cost of k-means clustering can be high, especially when frequent updates necessitate re-clustering.
@@
Hash-based Partitioning. Alternatively, some VDBs adopt hash-based partitioning, such as consistent or uniform hashing [weaviate, vespa, guo2022manu]. Specifically, the hashing partitioning strategy uses a hash function to map data to different partitions. The hash value of each data point determines which partition it belongs to. In VDBs, the hash value is typically calculated based on certain features of the vector (for example, values of specific dimensions or the entire vector). The hash function can evenly distribute data across partitions, preventing any single partition from storing too much data. However, when the data distribution is uneven, it may lead to some partitions becoming overloaded. Additionally, when the node number changes, hash-based partitioning may require the redistribution of the large dataset, which can incur significant overhead.
2.3 Caching
Caching plays a central role in optimizing the efficiency of data-intensive systems. In the context of VDBs, caching helps reduce latency and improve retrieval speed by temporarily storing vectors that are frequently accessed or recently used. While traditional databases often rely on in-memory key-value stores like Redis to cache structured data using predefined query keys, this approach does not directly translate to the vector space. The reason lies in the nature of the data: vectors are high-dimensional and continuous, making exact matches across queries rare. As VDBs increasingly serve workloads such as similarity search or embedding retrieval, the need for tailored caching strategies becomes clear. General caching algorithms—originally designed for web content or structured data—must be revisited in light of the unique access patterns and data characteristics found in vector systems. In this section, we discuss four widely used caching strategies—first-in first-out (FIFO), least recently used (LRU), most recently used (MRU), and least frequently used (LFU)—as well as partitioned caching in VDB design.
First-In First-Out (FIFO). FIFO is one of the simplest cache eviction strategies. It works by removing the oldest items in the cache once space runs out [grund2009abstract, grund2010precise]. Internally, it maintains a queue where new items are added to the back and removals occur from the front [mattson1970evaluation]. FIFO performs well in environments where data arrives in regular intervals and older entries quickly lose relevance—such as time-series data from IoT sensors. Its appeal lies in its constant-time performance and ease of implementation. However, because it ignores how often items are accessed, it may evict frequently used vectors prematurely.
Least Recently Used (LRU). The LRU algorithm prioritizes keeping recently accessed items in memory. When the cache fills up, it evicts the vector that hasn’t been used for the longest time, thereby improving the likelihood of cache hits [mattson1970evaluation]. This strategy aligns with many real-world patterns—such as search engines or recommendation systems—where recent queries are likely to be repeated. Systems like Redis have incorporated LRU for managing in-memory vectors. That said, LRU can struggle in workloads where access patterns are periodic or long-tailed, since it doesn’t track long-term popularity.
Most Recently Used (MRU). MRU takes the opposite stance: it evicts the most recently accessed item [gu2011theory]. This approach can be useful in scenarios where users rarely re-access the same vector right after retrieval, such as in certain data processing pipelines or streaming tasks. However, in most applications where temporal locality is strong, MRU tends to underperform, as it may discard items that will be queried again soon.